OpenStudy (anonymous):

Subtract (9X^4+5X^3-8)-(8X^2-2X+8)

OpenStudy (callisto):

Subtract the terms with same variable of the same power, so (9X^4+5X^3-8)-(8X^2-2X+8) = (9X^4)+(5X^3)-(8X^2)-(-2X)-8 -(+8) Can you work from here?

OpenStudy (anonymous):

Is this right 9x^4+5x^3-8x^2+2

OpenStudy (callisto):

Nope 9x^4+5x^3-8x^2 => correct +2 => incorreect and you've missed a term

OpenStudy (anonymous):

9x^4+5x^3-8x^2+2x

OpenStudy (callisto):

You've missed the constant term.

OpenStudy (callisto):

-8 -(+8) = -8-8, what would you get from here?

OpenStudy (anonymous):

-16

OpenStudy (anonymous):

9x^4+5x^3-8x^2+2x-16

OpenStudy (callisto):

Yes :)

OpenStudy (anonymous):

Thanx

OpenStudy (callisto):

You're welcome :)